OpenStudy (anonymous):

Laurie and Donnie are playing a marble game. According to their rules, 3 points are awarded for each red marble collected and 2 points are awarded for each green marble collected. Part A: Write an expression that will determine the score if they had r, red marbles and g, green marbles. Part B: How many points would be given for 5 red and 6 green marbles?

OpenStudy (anonymous):

scores=3*r+2*g ..........part A

OpenStudy (anonymous):

part B scores =3*5+2*6 score=15+12 s=27